OpenStudy (anonymous):

1. Complete the table below to solve the equation 2.5x – 10.5 = 64(0.5x). x f(x) = 2.5x – 10.5 g(x) = 64(0.5x) 2 3 4 5 6

OpenStudy (anonymous):

i need help with the 64(.5^x) parts

OpenStudy (displayerror):

Is the 'x' in the second part an exponent, or are you multiplying 0.5*x? If it is an exponent, it is just telling you that the base (in this case, 0.5) is being multiplied by itself some number (x) times. Thus, 0.5^2 (0.5 squared) = 0.5 * 0.5 = 0.25, 0.5^3 = 0.5 * 0.5 * 0.5 = 0.125